# Identification of Semiparametric Panel Multinomial

### v2

@@ 6.1 Data and Methodology
Specifically, we nonparametrically regress $\left(s_{ijt}-s_{ijs}\right)$ on the second-order polynomial basis functions of $\mathbf{X}_{i,ts}$ with $\ell_{1}$-regularization and 10-fold cross validation, and obtain an estimator $\hat{\gamma}_{j}$ of $\gamma_{j}\left(\overline{{\bf X}},\underline{{\bf X}}\right):=\mathbb{E}\left[\left.s_{ijt}-s_{ijs}\right|{\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right)\right]$. Then, we plug $\hat{\gamma}$ into our second-stage algorithm and compute the (approximate) argmin set $\hat{B}_{\hat{c}}$.
6.2 Results and Discussion
Table 4: Empirical Application: Estimation Results
$\hat{\beta}_{\hat{c}=0}^{m}$ $\left[\hat{\beta}^{l},\hat{\beta}^{u}\right]_{\hat{c}=0}$ $\hat{\beta}_{\hat{c}=.14}^{m}$ $\left[\hat{\beta}^{l},\hat{\beta}^{u}\right]_{\hat{c}=.14}$
@@ 6.2 Results and Discussion
[…]
The most interesting result is the positive estimated coefficient on the interaction term $\text{Price}{}_{ijt}$$\times$ $\text{Promo}_{ijt}$. An intuitive explanation for the positive sign is that by displaying certain products in front rows, consumers no longer see their price tags adjacent to those of their competitors, and thus become less price-sensitive for these specially promoted products.
Table 5: Empirical Illustration: Comparison of Results
$\hat{\beta}_{\hat{c}=.14}^{m}$ $\hat{\beta}^{CyclicMono}$ $\hat{\beta}^{OLS}$ $\hat{\beta}^{OLS-FE}$ $\hat{\beta}^{MLogit-FE}$
Priceijt -.9530 -.$\text{-}.3781$$0240$$\text{-}.3803$.$\text{-}.8511$
$\text{Promo}_{ijt}$1700 -.$\text{-}.0567$$5760$$5978$4589.4589.4589.4589
$\text{Price}{}_{ijt}$$\times$ $\text{Promo}_{ijt}$2324 .$9240$$\text{-}.8171$.$\text{-}.7057$.$\text{-}.2552$
Furthermore, we compare our $\hat{\beta}^{m}$ with the estimates obtained through four other methods, i.e. Cyclic Monotonicity (CM) based on Shi, Shum, and Song (2018) [^13] We used 2-week cycles for all available weeks in the data for the CM method., OLS, OLS with scalar-valued fixed effects (OLS-FE) and the multinomial logit with fixed effects (MLogit-FE). Results (normalized to $\mathbb{\mathbb{S}}^{D-1}$) are summarized in Table 5.
[…]
We regard the contrast between our result and the results obtained in these alternative methods as an empirical illustration that by accommodating more flexible forms of unobserved heterogeneity through the arbitrary dimensional fixed effects that are allowed to enter into consumers’ utility functions in an additively nonseparable way, our method is able to produce economically more reasonable results.
6.3 A Possible Explanation via Monte-Carlo Simulations
We provide a possible explanation here to the empirical findings in Table 5 via simulations. Recall that “Promo” captures whether a product gains increased exposure by being highlighted by stores. We argue that the negative estimated coefficients obtained in traditional methods in Table 5 for $\text{Price}{}_{ijt}$$\times$ $\text{Promo}_{ijt}$ may be caused by a positive correlation between display promotion and an unobserved index sensitivity term.
Specifically, suppose the utility function can be written as
where $X_{ijt}$ contains Price, Promo, and Price×$\times$×Promo, $A_{ij}$ is the $ij-$specific fixed effect which may capture index sensitivity (which can be thought as inversely related to unobserved brand loyalty), and $\epsilon_{ijt}$ is the exogenous random shock. Suppose $A_{ij}$ and Promoijt are positively correlated, which is reasonable because marketing managers with their expertise are more likely to promote products to which consumers are more price and promotion sensitive. Thus, traditional estimation methods that base on linearity would be unable to detect such pattern and wrongly attribute the effect on price elasticities from $A_{ij}$ to Promo.
[…]
We run 1,000 simulations for each of the five methods in Table 5 to estimate $\beta_{0}$. To replicate the data structure of the empirical exercise, we set $N=205$, $D=3$, $J=4$, and $T=10$. We report in Table 6 the percentage of simulations that the corresponding method is able to generate correct signs for all coordinates of $X_{ijt}$.
Table 6: Percentage of Correct Signs of Estimated Coefficients
$\alpha$ $\hat{\beta}^{m}$ $\hat{\beta}^{CyclicMono}$ $\hat{\beta}^{OLS}$ $\hat{\beta}^{OLS-FE}$ $\hat{\beta}^{MLogit-FE}$
 .15 92.40% 0.00% 0.00% 0.00% 27.40%
 .30 86.00% 0.00% 0.00% 0.00% 0.20%
 .50 74.10% 0.00% 0.00% 0.00% 0.00%
[…]
Intuitively, since products with larger $A_{ij}$ are more likely to be promoted $\left(X_{ijt,2}=1\right)$ by the selection of marketing managers, the average effective price sensitivity of promoted products tends to be larger than those products not promoted. This drives those estimators that ignore such confounding selection effects to produce a negative coefficient on the interaction term in Table 5. In contrast, our method handles such non-additive dependence between observable characteristics and unobserved fixed effects well, illustrating its robustness in these models.
Acknowledgments
We are grateful to Xiaohong Chen, Peter Phillips, and Phil Haile for their invaluable advice and encouragement. We thank Don Andrews, Isaiah Andrews, Tim Armstrong, Tim Christensen, Ben Connault, Francis Diebold, Bo Honoré, Joel Horowitz, Yuichi Kitamura, Patrick Kline, Lixiong Li, Yuan Liao, Charles Manski, Aviv Nevo, Matt Seo, Xiaoxia Shi, Frank Schorfheide, Elie Tamer, Ed Vytlacil, Rui Wang, Sheng Xu and seminar participants at Georgetown, UCSD, Berkeley, UCL, Northwestern, UW-Madison, UPenn, NYU, Princeton, HKU, CUHK, SMU, LSE, KU Leuven, Sciences Po, PSU, Harvard-MIT, Microeconometrics Class of 2019 Conference (Duke) and 2020 Winter Meeting of the Econometric Society for helpful comments. We thank Chuyue Tian for excellent research assistance.
A   Proof of Theorem 2
@@
Recalling that $\cdot v_{k}\left(\overline{{\bf X}}-\underline{{\bf X}}\right)=\overline{X}_{k}-\underline{X}_{k}/\left\|\overline{X}_{k}-\underline{X}_{k}\right\|$ whenever $\overline{X}_{k}\neq\underline{X}_{k}$ while $v_{k}\left(\overline{{\bf X}}-\underline{{\bf X}}\right)=0$ when $\overline{X}_{k}=\underline{X}_{k}$, we have
$G\left(\gamma_{j,t,s}\left({\bf X}_{i,ts}\right)\right)\lambda_{j}\left({\bf X}_{i,ts};\beta\right)=$ $G\left(\gamma_{j,t,s}\left({\bf X}_{i,ts}\right)\right)\prod_{k=1}^{J}\mathbf{\mathbbm{1}}\left\{\left(-1\right)^{\mathbf{\mathbbm{1}}\left\{k=j\right\}}\left(X_{ikt}-X_{iks}\right)^{{}^{\prime}}\beta\geq 0\right\}$
$=$ $G\left(\gamma_{j,t,s}\left({\bf X}_{i,ts}\right)\right)\prod_{k=1}^{J}\mathbf{\mathbbm{1}}\left\{\left(-1\right)^{\mathbf{\mathbbm{1}}\left\{k=j\right\}}v_{k}\left({\bf X}_{it}-{\bf X}_{is}\right)^{{}^{\prime}}\beta\geq 0\right\}$
which is continuous in $\beta$ with probability one, since $v_{k}\left({\bf X}_{it}-{\bf X}_{is}\right)$ has no mass point except possibly at ${\bf 0}$, in which case the indicator degenerates to a constant over $\beta\in\mathbb{\mathbb{S}}^{d-1}$. Since ${\bf X}_{i,ts}$ is i.i.d. across $i$, $\mathbb{\mathbb{S}}^{d-1}$ is compact, and the indicator function is bounded, all conditions for Lemma 2.4 in Newey and McFadden (1994) are satisfied, by which we conclude that $Q=\sum_{j,t,s}Q_{j,t,s}$ is continuous on $\mathbb{\mathbb{S}}^{d-1}$. ∎
@@
Under Assumptions 1, 5 and 6, $\sup_{\beta\in\mathbb{\mathbb{S}}^{d-1}}\left|\hat{Q}\left(\beta\right)-Q\left(\beta\right)\right|=O_{p}\left(c_{N}\right).$

### v3

@@ 6.2 Results and Discussion
We view the contrast between our findings and those from alternative methods as empirical evidence that, by accommodating more flexible forms of unobserved heterogeneity—via high-dimensional fixed effects that enter consumers’ utility functions in an additively nonseparable manner—our approach yields more economically plausible results.
6.3 A Possible Explanation via Monte-Carlo Simulations
In this subsection, we provide a possible explanation for the empirical findings reported in Table 4 through simulation analysis. Recall that “Promo” indicates whether a product receives increased in-store exposure by being highlighted by the store. We argue that the negative estimates on $\text{Price}{}_{ijt}$ $\times$ $\text{Promo}_{ijt}$ reported for traditional methods in Table 4 likely arise from a positive correlation between display promotions and an unobserved index of price sensitivity.
Specifically, suppose the utility function is
@@ 6.3 A Possible Explanation via Monte-Carlo Simulations
To provide some numerical evidence of the claim, we run the following Monte Carlo simulation. We set $\beta_{0}=\left(-4,2,2\right)^{{}^{\prime}}$, $Z_{ij}\sim\mathcal{U}\left[0,1\right]$, $A_{ij}=Z_{ij}+1$, and $\epsilon_{ijt}\sim TIEV\left(0,1\right)$. For the $X_{ijt}$ vector, we draw $X_{ijt,1}\sim\mathcal{U}\left[0,4\right]\text{ and }W_{ijt}\sim\mathcal{U}\left[0,1\right],$ and let $X_{ijt,2}=\left(1-\alpha\right)\times W_{ijt}+\alpha\times Z_{ij}$ and $X_{ijt,3}=X_{ijt,1}\times X_{ijt,2}$. We emphasize that $X_{ijt,2}$ (Promo) is positively correlated with $A_{ij}$ through $Z_{ij}$, with $\alpha$ measuring the strength of the correlation. We consider three values of $\alpha$: $0.15,\ 0.3,\text{ and }0.5$.
We run 1,000 simulations for each of the five methods in Table 4 to estimate $\beta_{0}$. To replicate the data structure of the empirical exercise, we set $N=205$, $D=3$, $J=4$, and $T=10$. We report in Table 5 the percentage of simulations that the corresponding method produces correct signs for all coordinates of $X_{ijt}$.
Table 5: Percentage of Correct Signs of Estimated Coefficients
$\alpha$ $\hat{\beta}^{m}$ $\hat{\beta}^{CyclicMono}$ $\hat{\beta}^{OLS}$ $\hat{\beta}^{OLS-FE}$ $\hat{\beta}^{MLogit-FE}$ $\hat{\beta}^{RCLM}$
 0.15 91.50% 0.00% 0.00% 0.00% 28.00% 0.00%
 0.30 85.90% 0.00% 0.00% 0.00% 0.20% 0.00%
 0.50 74.20% 0.00% 0.00% 0.00% 0.00% 0.00%
The percentages of simulations where our proposed method produces correct signs for all coordinates of $X_{ijt}$ for $\alpha=0.15,\ 0.3,$ and $0.5$ are 91.50%, 85.90%, and 74.20%, respectively. The accuracy of the estimator is negatively affected by the correlation between $X_{ijt,2}$ (Promo) and $A_{ij}$ (multiplicative fixed effect). In contrast, none of the other methods in Table 5 generates correct signs as ours does. The alternative models, owing to their additively separable structure, [^22] We note that the CM method requires $A_{ij}$ entering the utility function linearly, which is violated in (20). may overlook the positive dependence between Promo and the multiplicative fixed effect $A_{ij}$, which can bias the resulting estimates. [^23] Notably, RCLM produces “wrong signs” in this simulation exercise, even though it yields the expected signs in the empirical application in Section 6.2. A plausible interpretation is that, while RCLM is more flexible than, for example, MLogit-FE, the unknown selection effect in this dataset may be insufficiently strong to cause RCLM to fail, yet strong enough for MLogit-FE and related methods to do so. This observation highlights the potential value of our method as a robustness-check tool.
Intuitively, since products with larger $A_{ij}$ are more likely to be promoted $\left(X_{ijt,2}=1\right)$ by the selection of marketing managers, the average effective price sensitivity of promoted products tends to be greater in magnitude than that of non-promoted products. This drives those estimators that ignore such selection effects to produce a negative coefficient on the interaction term. In contrast, our method handles such non-additive dependence between observable characteristics and unobserved fixed effects well, illustrating its robustness in these models.
7 Conclusion
This paper develops a method for semiparametric identification and estimation in panel multinomial choice models that feature infinite-dimensional fixed effects and nonadditive utility, thereby accommodating rich forms of unobserved heterogeneity. We also introduce a general identification strategy based on multivariate monotonicity of parametric indices, applicable to a broad class of econometric models defined by the MISC conditions. In addition, we present a computational algorithm that leverages angle-space reparameterization and adaptive-grid search, which prove effective given the nonstandard criterion function implied by our identifying restrictions.
Future research could investigate how our approach might be applied and adapted to other specific microeconometric models within the MISC framework. Along this line, Gao, Li, and Xu (2023)—a companion paper to the present study—illustrates how the approach proposed here can be adapted to the context of dyadic network formation models. Gao and Wang (2025) propose a method for addressing endogeneity in discrete choice models under an adapted time-homogeneity condition. In ongoing work, we are investigating how to combine techniques developed in this line of work to analyze strategic network formation models with endogenous covariates.
Appendix A Proof of Theorem 2
Proof.
Let $\beta^{*}\in B_{0}\backslash\left\{\beta_{0}\right\}$. In the following, we condition on ${\bf x}_{st}$ and suppress ${\bf x}_{st}$ for notational simplicity. Write $p_{j\left(t\right)}:=\mathbb{P}\left(\left.y_{ijt}=1\right|{\bf x}_{st}\right)$. Let ${\cal J}_{+}^{*}:=\left\{j\in{\cal J}:\left(x_{s}-x_{t}\right)^{{}^{\prime}}\beta^{*}\geq 0\right\}$ and ${\cal J}_{-}^{*}:={\cal J}\backslash{\cal J}_{+}^{*}$. Then, we have
$\sum_{j\in{\cal J}_{+}}p_{j\left(s\right)}\geq\sum_{j\in{\cal J}_{+}}p_{j\left(t\right)}.$ (21)
Since $\sum_{j}p_{j\left(s\right)}=\sum_{j}p_{j\left(t\right)}=1$, we also have
$\sum_{j\notin{\cal J}_{+}}p_{j\left(s\right)}\leq\sum_{j\notin{\cal J}_{+}}p_{j\left(t\right)}.$ (22)
Without loss of generality, relabel products so that
${\cal J}_{+}^{*}=\left\{J,\ldots,j^{*}\right\},\quad{\cal J}_{-}^{*}=\left\{j^{*}-1,\ldots,1\right\},$
and
$p_{J\left(s\right)}-p_{J\left(t\right)}\geq\ldots\geq p_{j^{*}\left(s\right)}-p_{j^{*}\left(t\right)},$ (23)
and
$p_{j^{*}-1,\left(s\right)}-p_{j^{*}-1,\left(t\right)}\geq\ldots\geq p_{1,\left(s\right)}-p_{1,\left(t\right)}.$ (24)
In words, products are relabeled according to the following lexicographic ascending order:
  • Products in $\cal J$$_{+}$ receives larger labels than products in $\cal J$$_{-}$;
  • Within each of $\cal J$$_{+}$ and $\cal J$$_{-}$, products are further sorted so that $p_{js}-p_{jt}$ is ascending in (new) product label $j$.
Given (21) and (23), we must have
$\sum_{j=h}^{J}\left(p_{j\left(s\right)}-p_{j\left(t\right)}\right)\geq 0\quad\forall h\geq j^{*}.$ (25)
In the meanwhile, given (22) and (24), we also have
$\sum_{j=1}^{h}\left(p_{j\left(s\right)}-p_{j\left(t\right)}\right)\leq 0\quad\forall h\leq j^{*}-1.$
Again, since $\sum_{j}p_{j\left(s\right)}=\sum_{j}p_{j\left(t\right)}=1$, the above analysis implies that
$\left(1-\sum_{j=1}^{h}p_{j\left(s\right)}\right)-\left(1-\sum_{j=1}^{h}p_{j\left(t\right)}\right)\geq 0\quad\forall h\leq j^{*}-1,$
which is equivalent to
$\sum_{j=h+1}^{J}\left(p_{j\left(s\right)}-p_{j\left(t\right)}\right)\geq 0,\quad\forall h\leq j^{*}-1,$
which is further equivalent to
$\sum_{j=h}^{J}\left(p_{j\left(s\right)}-p_{j\left(t\right)}\right)\geq 0,\quad\forall h=1,\ldots,j^{*}.$ (26)
Combining (25) and (26) gives
$\sum_{j=h}^{J}\left(p_{j\left(s\right)}-p_{j\left(t\right)}\right)\geq 0\quad\forall h=1,\ldots,J.$ (27)
Following Pakes and Porter (2024), define the choice mapping $y\left({\bf x}_{s},a,\epsilon_{s},\beta\right)$ to the product label that is chosen according to model (1) at period $s$ with $\left({\bf x}_{s},a,\epsilon_{s},\beta\right)$, i.e.,
$y\left({\bf x}_{s},a,\epsilon_{s},\beta\right)=j\quad\Leftrightarrow\quad u\left(x_{js}^{{}^{\prime}}\beta,a_{j},\epsilon_{js}\right)>\max_{k\neq j}u\left(x_{ks}^{{}^{\prime}}\beta,a_{k},\epsilon_{ks}\right).$
Take $a$ to be any in-support value in $\cal A$}. Define
$R_{j;s}^{*}:=\left\{\tilde{\epsilon}_{s}:y\left({\bf x}_{s},a,\tilde{\epsilon}_{s},\beta^{*}\right)=j\right\}$
to be the values of $\epsilon_{s}^{*}$ that lead to product $j$ being chosen under $\left({\bf x}_{s},a,\beta^{*}\right)$ in period $s$.
Recall that $\left(x_{js}-x_{jt}\right)^{{}^{\prime}}\beta^{*}\geq 0$ for $j\in{\cal J}_{+}^{*}$ and $\left(x_{js}-x_{jt}\right)^{{}^{\prime}}\beta^{*}<0$ for $j\in{\cal J}\backslash{\cal J}_{+}^{*}$. Hence, by the monotonicity of $u$ in its first argument,
$\bigcup_{j\in{\cal J}_{+}^{*}}R_{j;t}^{*}\subseteq\bigcup_{j\in{\cal J}_{+}^{*}}R_{j;s}^{*},$ (28)
and hence
$R_{j;s}^{*}\cap R_{k;t}^{*}=\emptyset\quad\forall j\in{\cal J}_{-}^{*},\ k\in{\cal J}_{+}^{*}.$
Let $R_{j,k}^{*}:=R_{j;s}^{*}\cap R_{k;t}^{*}$ and
$R_{j,k,h,l}^{*}:=R_{j,k}^{*}\times R_{h,l}^{*}:=\left\{\left(\tilde{\epsilon}_{s},\tilde{\epsilon}_{t}\right):\tilde{\epsilon}_{s}\in R_{j,k}^{*},\tilde{\epsilon}_{t}\in R_{h,l}^{*}\right\}.$
Whenever $R_{j,k}^{*}\neq\emptyset$, pick any single point $r_{j,k}^{*}\in R_{j,k}^{*}$.
We specify the joint distribution of $\left(\epsilon_{s}^{*},\epsilon_{t}^{*}\right)$ as a discrete distribution over points $\left(r_{j,k}^{*},r_{h,l}^{*}\right)$. We write
$q_{j,k,h,l}^{*}:=\mathbb{P}\left(\left.\left(\epsilon_{s}^{*},\epsilon_{t}^{*}\right)=\left(r_{j,k}^{*},r_{h,l}^{*}\right)\right|{\bf x}_{s},{\bf x}_{t}\right)\equiv\mathbb{P}\left(\left.\left(\epsilon_{s}^{*},\epsilon_{t}^{*}\right)\in R_{j,k,h,l}^{*}\right|{\bf x}_{s},{\bf x}_{t}\right).$
Hence, a vector $\overline{q}^{*}=\left(q_{j,k,h,l}^{*}\right)$ in the unit simplex defines a joint distribution of $\left(\epsilon_{s}^{*},\epsilon_{t}^{*}\right)$ .
By (28), we have
$q_{j,k,h,l}^{*}=0,\ \forall\left(j,k\right)\text{ or }\left(h,l\right)\in{\cal J}_{-}^{*}\times{\cal J}_{+}^{*}.$ (29)
We set
$q_{j,k,h,l}^{*}=0,\ \forall j<k\text{ or }h<l.$ (30)
Note that $\left(j,k\right)\in{\cal J}_{-}^{*}\times{\cal J}_{+}^{*}$ implies $j<k$ by the relabeling. Hence, (30) sets a larger class of $q_{jkhl}$ to be zero than those in (29).
Given (30), to specify a joint distribution of $\left(\epsilon_{s}^{*},\epsilon_{t}^{*}\right)$, we only need to specify
$q^{*}:=\left(q_{j,k,h,l}^{*}\right)_{j\geq k,h\geq l},$ (31)
so that $\overline{q}^{*}:=\left(q^{*},\underline{q}^{*}\right)$ with
$\underline{q}^{*}:=\left(q_{j,k,h,l}^{*}\right)_{j<k\text{ or }h<l}={\bf 0}.$
First, $q^{*}$ needs to satisfy a homogeneity condition. Note that, given (30),
$\mathbb{P}\left(\left.\epsilon_{s}^{*}=r_{jk}\right|{\bf x}_{s},{\bf x}_{t}\right)$ $=\sum_{h,l}q_{j,k,h,l}^{*}=\sum_{h\geq l}q_{j,k,h,l}^{*}$
$\mathbb{P}\left(\left.\epsilon_{t}^{*}=r_{jk}\right|{\bf x}_{s},{\bf x}_{t}\right)$ $=\sum_{h,l}q_{h,l,j,k}^{*}=\sum_{h\geq l}q_{h,l,j,k}^{*}$
For $j<k$, conditional homogeneity is trivially satisfied, since
$\mathbb{P}\left(\left.\epsilon_{s}^{*}=r_{jk}\right|{\bf x}_{s},{\bf x}_{t}\right)=\mathbb{P}\left(\left.\epsilon_{t}^{*}=r_{jk}\right|{\bf x}_{s},{\bf x}_{t}\right)=0\text{ }\forall j<k.$
Hence, to satisfy homogeneity condition, we just need to ensure
$\sum_{h\geq l}q_{j,k,h,l}^{*}=\sum_{h\geq l}q_{h,l,j,k}^{*},\quad\forall j\geq k.$ (32)
Second, $q^{*}$ needs to produce choice probabilities that match the true probabilities. Let
$p_{jk}^{*}$ $:=\mathbb{P}\left(\left.y\left({\bf x}_{s},a,\epsilon_{s}^{*},\beta\right)=j,y\left({\bf x}_{t},a,\epsilon_{t}^{*},\beta\right)=k\right|{\bf x}_{s},{\bf x}_{t}\right)$
$=\mathbb{P}\left(\left.\left(\epsilon_{s}^{*},\epsilon_{t}^{*}\right)\in R_{j;s}^{*}\times R_{k;t}^{*}\right|{\bf x}_{s},{\bf x}_{t}\right).$
Then, the joint choice probabilities
$p_{jk}^{*}=\sum_{h=1}^{J}\sum_{l=1}^{J}q_{j,h,l,k}^{*}=\sum_{h\leq j}\sum_{l\geq k}q_{j,h,l,k}^{*}$
To match with true probabilities, we need
$\sum_{h\leq j}\sum_{l\geq k}q_{j,h,l,k}^{*}=p_{jk},\forall j,k=1,\ldots,J$ (33)
Given $p_{j\left(s\right)}=\sum_{k}p_{jk}$ and $p_{j\left(t\right)}=\sum_{k}p_{kj}$, we can now translate condition (27) about $p_{j\left(s\right)}-p_{j\left(t\right)}$ into a restriction about $\left(p_{jk}\right)$:
$\sum_{j=h}^{J}\sum_{k}p_{jk}\geq\sum_{j=h}^{J}\sum_{k}p_{kj},\quad\forall h=1,\ldots,J.$ (34)
To summarize, given (27), a condition on $p_{jk}$ that we know to hold, we need to show that there exists a $q^{*}:=\left(q_{j,k,h,l}^{*}\right)_{j\geq k,h\geq l}$ such that (32) and (33) both hold.
The above, however, is exactly the same as the one solved in Pakes and Porter (2024), as summarized by the following lemma. ∎
Lemma 1 (Nonnegative Solvability of Linear System in Pakes and Porter (2024)).
Let $q^{*}\equiv\left(q_{jkhl}^{*}\right)$ be a $\left(\frac{1}{2}J\left(J+1\right)\right)^{2}$-dimensional vector indexed by $j,k,h,l\in\left\{1,\ldots,J\right\}$ such that
$j\geq k\quad\text{and}\quad h\geq l.$
Let $p\equiv\left(p_{jk}\right)$ be a $J^{2}$-dimensional vector indexed by $j,k\in\left\{1,\ldots,J\right\}$ s.t. $p\in\Delta^{J^{2}-1}$, i.e., $p$ is a (discrete) probability distribution over $J^{2}$ points defined by $jk$.
Suppose that
$\sum_{j=h}^{J}\sum_{k}p_{jk}\geq\sum_{j=h}^{J}\sum_{k}p_{kj},\quad\forall h=1,\ldots,J.$ (35)
Then, there exists a nonnegative $q^{*}\geq 0$ that solves the following joint system of equations:
$\sum_{h\leq j}\sum_{l\geq k}q_{j,h,l,k}^{*}=p_{jk},$ $\quad\forall j,k\in\left\{1,\ldots,J\right\}.$ (36)
@@ Appendix A Proof of Theorem 2
Lemma 1 above summarizes the part of Pakes and Porter (2024)’s proof of their Theorem 2 that is useful to our setup. Specifically, the inequality constraints in (35) are the same as those in equation (S5) of Pakes and Porter (2024); equations in (36) are the same as those in (S3) of Pakes and Porter (2024); equations in (37) are the same as those in (S4) of Pakes and Porter (2024). The proof in Pakes and Porter (2024) after their equation (S5) then establishes the conclusion of Lemma 1.
Appendix B Proof of Theorem 3
@@
Recalling that $v_{k}\left(\overline{{\bf X}}-\underline{{\bf X}}\right)=\overline{X}_{k}-\underline{X}_{k}/\left\|\overline{X}_{k}-\underline{X}_{k}\right\|$ whenever $\overline{X}_{k}\neq\underline{X}_{k}$ while $v_{k}\left(\overline{{\bf X}}-\underline{{\bf X}}\right)=0$ when $\overline{X}_{k}=\underline{X}_{k}$, we have
$G\left(\gamma_{\tilde{{\cal J}},t,s}\left({\bf X}_{i,ts}\right)\right)\lambda_{j}\left({\bf X}_{i,ts};\beta\right)=$ $\ G\left(\gamma_{\tilde{{\cal J}},t,s}\left({\bf X}_{i,ts}\right)\right)\prod_{k=1}^{J}\mathbf{\mathbbm{1}}\left\{\left(-1\right)^{\mathbf{\mathbbm{1}}\left\{k\in\tilde{{\cal J}}\right\}}\left(X_{ikt}-X_{iks}\right)^{{}^{\prime}}\beta\geq 0\right\}$
$=$ $\ G\left(\gamma_{\tilde{{\cal J}},t,s}\left({\bf X}_{i,ts}\right)\right)\prod_{k=1}^{J}\mathbf{\mathbbm{1}}\left\{\left(-1\right)^{\mathbf{\mathbbm{1}}\left\{k\in\tilde{{\cal J}}\right\}}v_{k}\left({\bf X}_{it}-{\bf X}_{is}\right)^{{}^{\prime}}\beta\geq 0\right\},$
which is continuous in $\beta$ with probability one, since $v_{k}\left({\bf X}_{it}-{\bf X}_{is}\right)$ has no mass point except possibly at $\bf 0$}, in which case the indicator degenerates to a constant over $\beta\in\mathbb{\mathbb{S}}^{d-1}$. Since ${\bf X}_{i,ts}$ is i.i.d. across $i$, $\mathbb{\mathbb{S}}^{d-1}$ is compact, and the indicator function is bounded, all conditions for Lemma 2.4 in Newey and McFadden (1994) are satisfied, by which we conclude that $Q=\sum_{\tilde{{\cal J}},t,s}Q_{\tilde{{\cal J}},t,s}$ is continuous on $\mathbb{\mathbb{S}}^{d-1}$. ∎
@@
Under Assumptions 1, 5, and 6, $\sup_{\beta\in\mathbb{\mathbb{S}}^{d-1}}\left|\hat{Q}\left(\beta\right)-Q\left(\beta\right)\right|=O_{p}\left(c_{N}\right).$